OpenStudy (jordan):

solve: (x+3)squared equals36

OpenStudy (watchmath):

\((x+3)=36\) \(x+3=\pm 6\) \(x=6-3=3\) or \(x=-6-3=-9\) So \(x=3\) or \(x=-9\)